# Xinxu Wei1       Kanhao Zhao1       Yong Jiao1       Lifang He1       Yu Zhang [^1] footnotemark: 1

### v1

@@ Abstract
As large language models (LLMs) continue to revolutionize AI research, there is a growing interest in building large-scale brain foundation models to advance neuroscience. While most existing brain foundation models are pre-trained on time-series signals or region-of-interest (ROI) features, we propose a novel graph-based pre-training paradigm for constructing a brain graph foundation model. In this paper, we introduce the Brain Graph Foundation Model, termed BrainGFM, a unified framework that leverages graph contrastive learning and graph masked autoencoders for large-scale fMRI-based pre-training. BrainGFM is pre-trained on a diverse mixture of brain atlases with varying parcellations, significantly expanding the pre-training corpus and enhancing the model’s ability to generalize across heterogeneous fMRI-derived brain representations. To support efficient and versatile downstream transfer, we integrate both graph prompts and language prompts into the model design, enabling BrainGFM to flexibly adapt to a wide range of atlases, neurological and psychiatric disorders, and task settings. Furthermore, we employ meta-learning to optimize the graph prompts, facilitating strong generalization to previously unseen disorders under both few-shot and zero-shot learning conditions via language-guided prompting. BrainGFM is pre-trained on 27 neuroimaging datasets spanning 25 common neurological and psychiatric disorders, encompassing 2 types of brain atlases (functional and anatomical) across 8 widely-used parcellations, and covering over 25,000 subjects, 60,000 fMRI scans, and a total of 400,000 graph samples aggregated across all atlases and parcellations. The code is available at https://github.com/weixinxu666/BrainGFM.
@@ 1 Introduction
With the rise of large language models (LLMs) Achiam et al. (2023), large-scale pre-trained foundation models (FMs) have been proposed across various domains, including computer vision Touvron et al. (2023), natural language processing Achiam et al. (2023), and data mining Xia et al. (2024). Recently, the field of neuroscience has also begun to witness the emergence of brain foundation models. As a widely used data modality in neuroscience, functional magnetic resonance imaging (fMRI) Markiewicz et al. (2021); Bycroft et al. (2018) plays a crucial role in understanding brain function and dysfunction. Developing a fMRI-based brain foundation model is of great importance for advancing neuroscience and its translational research. Due to the high complexity and cost of fMRI data acquisition Van Essen et al. (2012); Cui et al. (2022), coupled with strong heterogeneity and substantial inter-subject variability, most existing traditional deep learning-based fMRI models Wei et al. (2025); Kan et al. (2022); Jiao et al. (2025) are trained on relatively small datasets. Consequently, these models are typically tailored to specific tasks, disorders, or cohorts, resulting in limited generalizability, poor flexibility, and weak transferability to unseen tasks, datasets or disorder conditions. Furthermore, training on small-scale datasets often leads to under-fitting, ultimately compromising model performance and reliability. These issues have become common limitations of traditional deep learning models for fMRI. However, they can be effectively addressed by building fMRI brain FMs. Since FMs are typically pre-trained on large-scale datasets with rich diversity Touvron et al. (2023); Xia et al. (2024), spanning various data types and knowledge representations within the neuroscience domain, the resulting models exhibit strong generalization and broad applicability.
@@
Previous fMRI FMs have uniformly adopted Transformer-based architectures and were exclusively pre-trained on either time-series data or ROI-level features, resulting in two main categories: time-series-based Caro et al. (2023); Thomas et al. (2022) and ROI-based Yang et al. (2024); Hu et al. (2024); Dong et al. (2024) fMRI brain FMs. However, building brain FMs faces several critical challenges that many previous approaches have either overlooked or failed to effectively address. (1) Data Availability & Heterogeneity. fMRI data are difficult and costly to collect and pre-process Poldrack and Gorgolewski (2017), yet pre-training FMs typically requires large-scale datasets. Existing fMRI datasets are not only limited in quantity but also exhibit substantial heterogeneity across sources. Effectively leveraging and integrating these heterogeneous datasets is thus a fundamental challenge. Many existing brain FMs have constructed relatively large-scale fMRI pre-training datasets, but these are typically based on a single brain parcellation or atlas Thomas et al. (2022). This overlooks the fact that integrating multiple parcellation templates can not only expand the scale of available fMRI data but also provide diverse and even complementary features across different brain parcellations Hermosillo et al. (2024). (2) Pre-Training Computational Cost. The computational cost of pre-training brain FMs typically depends on the form of the fMRI data and the chosen pre-training strategy. Time-series-based brain FMs are pre-trained directly on raw fMRI time series, resulting in high computational demands with masked modeling pre-training paradigm. While ROI-based brain FMs are more efficient, they often neglect inter-regional connectivity, leading to suboptimal performance on various downstream tasks. Striking a balance between computational efficiency and modeling effectiveness remains a pressing issue. (3) Adaptability and Generalization for Few/Zero-Shot Transfer. Pre-trained brain FMs needs to be fine-tuned to various downstream tasks, datasets, atlases and disorders. However, full-parameter fine-tuning is often inefficient, requires large amounts of labeled data, and typically previous brain FMs Caro et al. (2023); Yang et al. (2024) support only one disorder or atlas during the downstream inference. In addition, in many real-world scenarios, downstream tasks may involve new atlases, datasets and disorders unseen during pre-training, with very limited (few-shot) or even no labeled data available (zero-shot). Adapting FMs to such few-shot or zero-shot settings poses a significant yet highly valuable challenge. Most existing brain FMs Caro et al. (2023); Thomas et al. (2022); Dong et al. (2024) have not considered few-shot or zero-shot scenarios, which limits their generalizability and flexibility. These three challenges correspond to four essential aspects of pre-training brain FMs: data collection, model pre-training & fine-tuning, and downstream task adaptation.
@@
Contributions. To address the key challenges outlined above and overcome the limitations of prior work, we propose the Brain Graph Foundation Model, named BrainGFM, specifically designed for heterogeneous fMRI data, with a particular focus on graph-based modeling. We propose corresponding solutions within our model to enhance BrainGFM, enabling it to become a more powerful brain FM compared to previous approaches. (1) To enable effective pre-training of brain FMs, we construct a large-scale fMRI dataset comprising 27 widely used fMRI datasets. This collection includes over 25,000 subjects, 60,000 fMRI scans, and 25 common neurological and psychiatric disorders. Unlike previous brain foundation models, each fMRI sample in our dataset is processed using 2 different brain functional and anatomical atlases, including 8 parcellations with various resolutions and partitions, significantly increasing the scale and diversity of the data. This also allows the pre-trained model to capture complementary feature representations across multiple parcellations. (2) Prior brain FMs have predominantly relied on fMRI time series or ROI-level features for both pre-training and fine-tuning. In this work, we creatively introduce a graph-based backbone for building brain graph FMs. This approach offers the advantage of maintaining computational efficiency comparable to ROI-based FMs, while achieving performance on par with time-series-based FMs. (3) To enhance the generalizability and adaptability of the model, we discard conventional fine-tuning and introduce a graph prompt-tuning. Under the multi-task and multi-dataset training paradigm of meta-learning, this approach improves the model’s ability to perform few-shot adaptation across diverse tasks and datasets. In addition, we incorporate language prompt tokens, including atlas/parcellation tokens and task/disorder tokens, to guide the pre-trained BrainGFM in adapting to entirely unseen downstream datasets, atlases, tasks, and disorders in zero-shot settings.
@@ 2.1 Pre-Training Approaches for Brain Foundation Models Using fMRI
The emergence of large-scale foundation models, such as LLMs Achiam et al. (2023), has demonstrated strong potential across various domains. In neuroscience, recent efforts have introduced brain FMs (e.g., using fMRI), which can be broadly classified into time-series-based Dong et al. (2024); Caro et al. (2023); Thomas et al. (2022) and ROI-based models Yang et al. (2024); Hu et al. (2024), both primarily relying on generative pre-training using masked modeling. In contrast to these approaches, our work introduces the first graph-based fMRI foundation model, which leverages the brain’s topological structure through graph representations. We incorporate both graph generative pre-training Hou et al. (2022) and graph contrastive pre-training Qiu et al. (2020); Wei et al. (2024), unifying two major paradigms in graph representation learning.
@@ 2.2 Graph Pre-Training and Prompt Learning
Pre-training is a fundamental step in the development of foundation models, with most approaches categorized into contrastive-based and generative-based paradigms. While graph model pre-training differs from that in vision and language domains, it generally follows these two strategies. To facilitate zero-shot generalization, language prompts Achiam et al. (2023) have been widely used in NLP and vision, providing semantic guidance that enables pre-trained models to adapt to unseen tasks without parameter updates. In contrast, graph prompts have been proposed to address few-shot adaptation in graph neural networks. Inspired by prefix-tuning Li and Liang (2021), graph prompts Sun et al. (2023) introduce a small set of task-specific parameters that can be optimized efficiently while keeping the backbone frozen. This approach improves sample efficiency and reduces computational cost in adapting to new graph-based tasks with limited data.
@@ 2.3 Meta-Learning
Meta-learning Finn et al. (2017); Hospedales et al. (2021), also known as “learn to learn” aims to train models that can quickly adapt to new tasks using only a small number of labeled examples. It typically involves learning a good initialization or adaptation strategy by optimizing over a distribution of related tasks. Meta-learning has been widely adopted in few-shot learning scenarios and has shown strong potential for improving generalization across tasks and domains Sun et al. (2023). In our study, meta-learning is employed to train the graph prompt under the few-shot setting, enabling the unification and generalization across diverse brain atlases and neurological disorders.
@@ 3 Methodology
As illustrated in Figure 1, we propose BrainGFM, a graph-based paradigm that distinguishes itself from previous time series-based and ROI-based brain FMs. Our framework consists of four main stages: large-scale fMRI graph data collection and pre-processing, graph pre-training for building our brain graph foundation model, multi-task meta-learning optimization for few-shot learning, and graph/language prompt-tuning for zero-shot adaption.
@@ 3.1 Construction of Large-Scale fMRI Pre-Training Dataset
As shown in Figure 1(a), we curated a large-scale fMRI dataset by aggregating 27 widely used fMRI datasets from different sites and institutions, covering 25 common neurological and psychiatric disorders. Unlike existing brain FMs, our dataset incorporates fMRI data processed using 8 parcellations, including Schaefer100/Schaefer200/Schaefer300 Schaefer et al. (2018), AAL116/AAL3v1 Tzourio-Mazoyer et al. (2002), SHEN268 Shen et al. (2013), Power264 Power et al. (2011), and Gordon333 Gordon et al. (2016). For each subject, we extracted raw fMRI time series using these brain atlases and constructed fMRI brain graphs by computing and binarizing the Pearson correlation between time series among brain ROIs. Integrating multiple atlases allows us to expand the dataset to eight times the size of using a single parcellation, enabling more diverse representations and facilitating the learning of atlas-invariant brain patterns. The inclusion of multiple atlases not only increases the diversity and volume of the training data but also enables the model to learn parcellation-specific features, significantly enhancing the generalization and robustness of the pre-trained BrainGFM. Note that detailed information regarding the benchmark settings, including task types, dataset splits, neurological disorder categories, and atlas/parcellation choices, can be found in Appendix N and P.
@@ 3.2 Graph Pre-Training for Building BrainGFM
Motivation. The graph foundation model approaches the effectiveness of time-series-based foundation models, while matching the efficiency of ROI-based foundation models.
@@
We adopt a Graph Transformer Yun et al. (2019) as the backbone of our BrainGFM. As illustrated in Figure 1(b), we first transform the input fMRI brain graphs and project them to obtain brain graph embeddings, where each token corresponds to a brain ROI. We apply Positional Encoding Tokens, denoted as [PE], to each brain ROI, enabling the model to perceive and learn the topological and spatial characteristics of each ROI. Unlike conventional positional encodings used in standard Transformer models Vaswani et al. (2017), graph-based positional encodings are inherently different, as they require encoding the relative positions between nodes in the graph structure. Compared to the commonly used Laplacian positional encoding Dwivedi et al. (2023) and node degree positional encoding You et al. (2019) in graph-based models, we adopt a more efficient alternative Random Walk Structural Encoding (RWSE) Dwivedi et al. (2021) as our positional encoding strategy. More details about PEs can be found in Appendix F. Furthermore, inspired by language models in NLP, we insert Atlas/Parcellation Tokens, denoted as [A/P], to the brain graph embeddings during pre-training, enabling the model to better distinguish and learn from different atlases and parcellations. Note that the construction of [A/P] tokens is described in detail in Section 3.4. Incorporating this token enables the model to capture parcellation-specific patterns, which is crucial as prior studies Hermosillo et al. (2024); Liu et al. (2023); Wu et al. (2025) show that different brain disorders are better represented by specific parcellations. For instance, MDD benefits from Schaefer200 or Power264 and ASD is better captured by Shen268 or Schaefer200. Embedding such parcellation-aware information helps improve model generalization across disorders and atlas configurations.
@@
We follow the graph pre-training paradigm to pre-train our BrainGFM. To fully leverage the potential of graph pre-training, we adopt two widely used pretext tasks in graph domain: graph contrastive learning (GCL) You et al. (2020) pre-training and graph masked autoencoders (GMAE) Hou et al. (2022) pre-training. For GCL pre-training, we apply graph augmentation to the fMRI brain graphs by randomly dropping nodes and edges to generate positive and negative pairs of queries and keys. The contrastive loss is then computed by contrasting the positive and negative graph pairs. For GMAE pre-training, we randomly mask nodes and edges in the input brain graphs to obtain masked brain graphs. These graphs are then passed through a graph autoencoder with an encoder-decoder architecture to reconstruct the masked nodes and edges, optimized using a mean squared error (MSE) loss. Note that both GCL and GMAE pre-training share the same encoder, which serves as the core of our BrainGFM, which enables the encoder of BrainGFM to benefit from both contrastive and generative paradigms, resulting in a more robust and well-pre-trained backbone. More details about these two graph pre-training methods can be found in Appendix J and K.

### v2

# Xinxu Weia    Kanhao Zhaoa    Yong Jiaoa    Lifang Hea    Yu Zhanga11footnotemark: 1

@@ Abstract
As large language models (LLMs) continue to revolutionize AI research, there is a growing interest in building large-scale brain foundation models to advance neuroscience. While most existing brain foundation models are pre-trained on time-series signals or connectome features, we propose a novel graph-based pre-training paradigm for constructing a brain graph foundation model. In this paper, we introduce the Brain Graph Foundation Model, termed BrainGFM, a unified framework that leverages graph contrastive learning and graph masked autoencoders for large-scale fMRI-based pre-training. BrainGFM is pre-trained on a diverse mixture of brain atlases with varying parcellations, significantly expanding the pre-training corpus and enhancing the model’s ability to generalize across heterogeneous fMRI-derived brain representations. To support efficient and versatile downstream transfer, we integrate both graph prompts and language prompts into the model design, enabling BrainGFM to flexibly adapt to a wide range of atlases, neurological and psychiatric disorders, and task settings. Furthermore, we employ meta-learning to optimize the graph prompts, facilitating strong generalization to previously unseen disorders under both few-shot and zero-shot learning conditions via language-guided prompting. BrainGFM is pre-trained on 27 neuroimaging datasets spanning 25 common neurological and psychiatric disorders, encompassing 2 types of brain atlases (functional and anatomical) across 8 widely-used parcellations, and covering over 25,000 subjects, 60,000 fMRI scans, and a total of 400,000 graph samples aggregated across all atlases and parcellations.
@@ 1 Introduction
With the rise of large language models (LLMs) (Achiam et al., 2023), large-scale pre-trained foundation models (FMs) have been proposed across various domains, including computer vision (Touvron et al., 2023), natural language processing (Achiam et al., 2023), and data mining (Xia et al., 2024). Recently, the field of neuroscience has also begun to witness the emergence of brain foundation models. As a widely used data modality in neuroscience, functional magnetic resonance imaging (fMRI) (Markiewicz et al., 2021; Bycroft et al., 2018) plays a crucial role in understanding brain function and dysfunction. Developing a fMRI-based brain foundation model is of great importance for advancing neuroscience and its translational research. Due to the high complexity and cost of fMRI data acquisition (Van Essen et al., 2012; Cui et al., 2022), coupled with strong heterogeneity and substantial inter-subject variability, most existing traditional deep learning-based fMRI models (Wei et al., 2025; Kan et al., 2022; Jiao et al., 2025) are trained on relatively small datasets. Consequently, these models are typically tailored to specific tasks, disorders, or cohorts, resulting in limited generalizability, poor flexibility, and weak transferability to unseen tasks, datasets or disorder conditions. Furthermore, training on small-scale datasets often leads to under-fitting, ultimately compromising model performance and reliability. These issues have become common limitations of traditional deep learning models for fMRI. However, they can be effectively addressed by building fMRI brain FMs. Since FMs are typically pre-trained on large-scale datasets with rich diversity (Touvron et al., 2023; Xia et al., 2024), spanning various data types and knowledge representations within the neuroscience domain, the resulting models exhibit strong generalization and broad applicability.
@@
Previous fMRI FMs have uniformly adopted Transformer-based architectures and were exclusively pre-trained on either time-series data or ROI-level features, resulting in two main categories: time-series-based (Caro et al., 2023; Thomas et al., 2022) and Connectome/FC-based (Yang et al., 2024; Hu et al., 2024; Dong et al., 2024a) fMRI brain FMs. However, building brain FMs faces several critical challenges that many previous approaches have either overlooked or failed to effectively address. (1) Data Availability & Heterogeneity. fMRI data are difficult and costly to collect and pre-process (Poldrack and Gorgolewski, 2017), yet pre-training FMs typically requires large-scale datasets. Existing fMRI datasets are not only limited in quantity but also exhibit substantial heterogeneity across sources. Effectively leveraging and integrating these heterogeneous datasets is thus a fundamental challenge. Many existing brain FMs have constructed relatively large-scale fMRI pre-training datasets, but these are typically based on a single brain parcellation or atlas (Thomas et al., 2022). This overlooks the fact that integrating multiple parcellation templates can not only expand the scale of available fMRI data but also provide diverse and even complementary features across different brain parcellations (Hermosillo et al., 2024). (2) Pre-Training Computational Cost. The computational cost of pre-training brain FMs typically depends on the form of the fMRI data and the chosen pre-training strategy. Time-series-based brain FMs are pre-trained directly on raw fMRI time series, resulting in high computational demands with masked modeling pre-training paradigm. While Connectome/FC-based brain FMs are more efficient, they often neglect inter-regional connectivity, leading to suboptimal performance on various downstream tasks. Striking a balance between computational efficiency and modeling effectiveness remains a pressing issue. (3) Adaptability and Generalization for Few/Zero-Shot Transfer. Pre-trained brain FMs need to be fine-tuned to various downstream tasks, datasets, atlases and disorders. However, full-parameter fine-tuning is often inefficient, requires large amounts of labeled data, and typically previous brain FMs (Caro et al., 2023; Yang et al., 2024) support only one disorder or atlas during the downstream inference. In addition, in many real-world scenarios, downstream tasks may involve new atlases, datasets and disorders unseen during pre-training, with very limited (few-shot) or even no labeled data available (zero-shot). Adapting FMs to such few-shot or zero-shot settings poses a significant yet highly valuable challenge. Most existing brain FMs (Caro et al., 2023; Thomas et al., 2022; Dong et al., 2024a) have not considered few-shot or zero-shot scenarios, which limits their generalizability and flexibility. These three challenges correspond to four essential aspects of pre-training brain FMs: data collection, model pre-training & fine-tuning, and downstream task adaptation.
@@
Contributions. To address the key challenges outlined above and overcome the limitations of prior work, we propose the Brain Graph Foundation Model, named BrainGFM, specifically designed for heterogeneous fMRI data, with a particular focus on graph-based modeling. We propose corresponding solutions within our model to enhance BrainGFM, enabling it to become a more powerful brain FM compared to previous approaches. (1) To enable effective pre-training of brain FMs, we construct a large-scale fMRI dataset comprising 27 widely used fMRI datasets. This collection includes over 25,000 subjects, 60,000 fMRI scans, and 25 common neurological and psychiatric disorders. Unlike previous brain foundation models, each fMRI sample in our dataset is processed using 2 different brain functional and anatomical atlases, including 8 parcellations with various resolutions and partitions, significantly increasing the scale and diversity of the data. This also allows the pre-trained model to capture complementary feature representations across multiple parcellations. (2) Prior brain FMs have predominantly relied on fMRI time series or ROI-level features for both pre-training and fine-tuning. In this work, we creatively introduce a graph-based backbone for building brain graph FMs. This approach offers the advantage of maintaining computational efficiency comparable to Connectome/FC-based FMs, while achieving performance on par with time-series-based FMs. (3) To enhance the generalizability and adaptability of the model, we discard conventional fine-tuning and introduce a graph prompt-tuning. Under the multi-task and multi-dataset training paradigm of meta-learning, this approach improves the model’s ability to perform few-shot adaptation across diverse tasks and datasets. In addition, we incorporate language prompt tokens, including atlas/parcellation tokens and task/disorder tokens, to guide the pre-trained BrainGFM in adapting to entirely unseen downstream datasets, atlases, tasks, and disorders in zero-shot settings.
@@ 2.1 Pre-Training Approaches for Brain Foundation Models Using fMRI
The emergence of large-scale foundation models, such as LLMs (Achiam et al., 2023), has demonstrated strong potential across various domains. In neuroscience, recent efforts have introduced brain FMs (e.g., using fMRI), which can be broadly classified into time-series-based (Dong et al., 2024a; Caro et al., 2023; Thomas et al., 2022) and Connectome/FC-based models (Yang et al., 2024; Hu et al., 2024), both primarily relying on generative pre-training using masked modeling. In contrast to these approaches, our work introduces the first graph-based fMRI foundation model, which leverages the brain’s topological structure through graph representations. We incorporate both graph generative pre-training (Hou et al., 2022) and graph contrastive pre-training (Qiu et al., 2020; Wei et al., 2024), unifying two major paradigms in graph representation learning.
@@ 2.2 Graph Pre-Training and Prompt Learning
Pre-training is a fundamental step in the development of foundation models, with most approaches categorized into contrastive-based and generative-based paradigms. While graph model pre-training differs from that in vision and language domains, it generally follows these two strategies. To facilitate zero-shot generalization, language prompts (Achiam et al., 2023) have been widely used in NLP and vision, providing semantic guidance that enables pre-trained models to adapt to unseen tasks without parameter updates. In contrast, graph prompts have been proposed to address few-shot adaptation in graph neural networks. Inspired by prefix-tuning (Li and Liang, 2021), graph prompts (Sun et al., 2023) introduce a small set of task-specific parameters that can be optimized efficiently while keeping the backbone frozen. This approach improves sample efficiency and reduces computational cost in adapting to new graph-based tasks with limited data. Recent studies have also introduced brain prompt-tuning methods (Xu et al., 2025; Dong et al., 2024b) to adapt pre-trained fMRI models.
@@ 2.3 Meta-Learning
Meta-learning (Finn et al., 2017; Hospedales et al., 2021), also known as “learn to learn” aims to train models that can quickly adapt to new tasks using only a small number of labeled examples. It typically involves learning a good initialization or adaptation strategy by optimizing over a distribution of related tasks. Meta-learning has been widely adopted in few-shot learning scenarios and has shown strong potential for improving generalization across tasks and domains (Sun et al., 2023). In our study, meta-learning is employed to train the graph prompt under the few-shot setting, enabling the unification and generalization across diverse brain atlases and neurological disorders.
@@ 3 Methodology
As illustrated in Figure 1, we propose BrainGFM, a graph-based paradigm that distinguishes itself from previous time series-based and Connectome/FC-based brain FMs. Our framework consists of four main stages: large-scale fMRI graph data collection and pre-processing, graph pre-training for building our brain graph foundation model, multi-task meta-learning optimization for few-shot learning, and graph/language prompt-tuning for zero-shot adaptation.
@@ 3.1 Construction of Large-Scale fMRI Pre-Training Dataset
As shown in Figure 1(a), we curated a large-scale fMRI dataset by aggregating 27 widely used fMRI datasets from different sites and institutions, covering 25 common neurological and psychiatric disorders. Unlike existing brain FMs, our dataset incorporates fMRI data processed using 8 parcellations, including Schaefer100/Schaefer200/Schaefer300 (Schaefer et al., 2018), AAL116/AAL3v1 (Tzourio-Mazoyer et al., 2002), SHEN268 (Shen et al., 2013), Power264 (Power et al., 2011), and Gordon333 (Gordon et al., 2016). For each subject, we extracted raw fMRI time series using these brain atlases and constructed fMRI brain graphs by computing and binarizing the Pearson correlation between time series among brain ROIs. Integrating multiple atlases allows us to expand the dataset to eight times the size of using a single parcellation, enabling more diverse representations and facilitating the learning of atlas-invariant brain patterns. The inclusion of multiple atlases not only increases the diversity and volume of the training data but also enables the model to learn parcellation-specific features, significantly enhancing the generalization and robustness of the pre-trained BrainGFM. Note that detailed information regarding the benchmark settings, including task types, dataset splits, neurological disorder categories, and atlas/parcellation choices, can be found in Appendix R and T.
@@ 3.2 Graph Pre-Training for Building BrainGFM
Motivation. The graph foundation model approaches the effectiveness of time-series-based foundation models, while matching the efficiency of Connectome/FC-based foundation models.
@@
We adopt a Graph Transformer (Yun et al., 2019) as the backbone of our BrainGFM. As illustrated in Figure 1(b), we first transform the input fMRI brain graphs and project them to obtain brain graph embeddings, where each token corresponds to a brain ROI. We apply Positional Encoding Tokens, denoted as [PE], to each brain ROI, enabling the model to perceive and learn the topological and spatial characteristics of each ROI. Unlike conventional positional encodings used in standard Transformer models (Vaswani et al., 2017), graph-based positional encodings are inherently different, as they require encoding the relative positions between nodes in the graph structure. Compared to the commonly used Laplacian positional encoding (Dwivedi et al., 2023) and node degree positional encoding (You et al., 2019) in graph-based models, we adopt a more efficient alternative Random Walk Structural Encoding (RWSE) (Dwivedi et al., 2021) as our positional encoding strategy. More details about PEs can be found in Appendix I. Furthermore, inspired by language models in NLP, we insert Atlas/Parcellation Tokens, denoted as [A/P], to the brain graph embeddings during pre-training, enabling the model to better distinguish and learn from different atlases and parcellations. Note that the construction of [A/P] tokens is described in detail in Section 3.4. Incorporating this token enables the model to capture parcellation-specific patterns, which is crucial as prior studies (Hermosillo et al., 2024; Liu et al., 2023; Wu et al., 2025) show that different brain disorders are better represented by specific parcellations. For instance, MDD benefits from Schaefer200 or Power264 and ASD is better captured by Shen268 or Schaefer200. Embedding such parcellation-aware information helps improve model generalization across disorders and atlas configurations.
@@
We follow the graph pre-training paradigm to pre-train our BrainGFM. To fully leverage the potential of graph pre-training, we adopt two widely used pretext tasks in graph domain: graph contrastive learning (GCL) (You et al., 2020) pre-training and graph masked autoencoders (GMAE) (Hou et al., 2022) pre-training. For GCL pre-training, we apply graph augmentation to the fMRI brain graphs by randomly dropping nodes and edges to generate positive and negative pairs of queries and keys. The contrastive loss is then computed by contrasting the positive and negative graph pairs. For GMAE pre-training, we randomly mask nodes and edges in the input brain graphs to obtain masked brain graphs. These graphs are then passed through a graph autoencoder with an encoder-decoder architecture to reconstruct the masked nodes and edges, optimized using a mean squared error (MSE) loss. Note that both GCL and GMAE pre-training share the same encoder, which serves as the core of our BrainGFM, which enables the encoder of BrainGFM to benefit from both contrastive and generative paradigms, resulting in a more robust and well-pre-trained backbone. More details about these two graph pre-training methods can be found in Appendix N and O.